Consider problem with no non-conservative forces. In this problem, if the 10 kg object in question starts from rest and speeds up to 10 m/s, what can we say about the change in the potential energy over the same time?

a) There is zero change in potential energy

b) The potential energy increases by roughly 50 J

c) The potential energy increases by roughly 500 J

d) The potential energy decreases by roughly 50 J

e) The potential energy decreases by roughly 500 J

Why is this answer correct
Answer #1

From law of conservation of energy

(K+U)i =(K+U)f

Ki +Ui =Kf +Uf

1/2*m*vi^2 +Ui =1/2*m*vf^2 +Uf

Uf - Ui =1/2*m*vi^2 - 1/2*m*vf^2

at rest,vi =0 m/s

Uf -Ui =0-1/2*10 kg*(10 m/s)^2

= -500 J

" -" means decreases.

answer is option (e)

=